Operations Report – March 20th

Operations Report 20-MAR-2020

SOL: n/a

Name of person filing report: Shannon Rupert

Non-nominal systems: None!

Notes on non-nominal systems: I find it just a little ironic that all systems are nominal right now.

Power system: SOLAR ONLY. Nominal. 99% at 8 pm last night, 77% at 7 am this morning.

Diesel Reading – 60%

Station Propane Reading – 68 %

Ethanol Free Gasoline – 5.5 gallons

Water (loft tank): full gallons

Water Meter: n/a units

Water (static tank) – 150 gallons

Static to Loft Pump used – yes

Water in GreenHab – 0 gallons

Water in ScienceDome: 0 gallons

Toilet tank emptied: no

Perseverance rover used: no

Hours: 152.5

Beginning charge: 100%

Ending charge: 100%

Currently charging: yes

Sojourner rover used: yes

Hours: 104.5

Beginning charge: 100%

Ending charge: 96 %

Currently charging: yes

Spirit rover: yes

Hours: 134.2

Beginning charge: 100%

Ending charge: 98 %

Currently charging: yes

Opportunity rover used: no

Hours: 70.3

Beginning charge: 100%

Ending charge: 100 %

Currently charging: yes

Curiosity rover used: yes

Hours: 139.6

Beginning charge: 100%

Ending charge: 100 %

Currently charging: yes

Notes on rovers: Nothing to report

ATV’s Used: No

Reason for use: n/a

Oil Added? no

ATV Fuel Used: None

Hours the ATVs were Used today: 0

Notes on ATVs: Nothing to report.

HabCar used and why, where? Just moved on station

CrewCar used and why, where? CrewCar is currently in Hanksville and will remain there for a while.

General notes and comments: Tunnel has been stripped of all tarps and tennis balls. They are stored in the RAM. The tennis ball caps worked perfectly all season.

The trash on the trailer has been taken to the dump. One load remains. This is a huge deal, as it has been sitting on the trailer for over a year. One of the tires on the trailer was low.

Summary of internet: Nominal.

Summary of suits and radios: All radios and spacesuits are no longer connected to power.

Summary of Hab operations: The lower deck lighting was replaced today with three 18, 000 LED lights. It’s delightfully daylight bright in there now! We also replaced the light in the EVA room with an LED daylight fixture. We should replace the remaining lower deck light

fixtures to complete the transformation.

Summary of GreenHab operations: Shutdown almost completed. I need to vacuum and bleach the floor and it will be complete

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Nothing to report.

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report.

Questions, concerns and requests to Mission Support: It’s been a good day and a lot got done!

Copyright © The Mars Society. All rights reserved. | Main Site